OpenStudy (anonymous):

Find the height of a rectangular prism if the surface area is 148 cm2, the width is 4 cm and the length is 5 cm.

OpenStudy (anonymous):

is that supposed to be 148cm^3? it's a three-dimensional object right?

OpenStudy (anonymous):

If so, 148 ----- = height (4*5)

OpenStudy (anonymous):

No. it's 148 cm2

Directrix (directrix):

@ehuman >>surface area is 148 cm2 AREA is expressed in square units.

OpenStudy (anonymous):

oh well now that i pay attention when i read it

Directrix (directrix):

SA of a rectangular solid is the perimeter of a base times the height of the solid PLUS the sum of the areas of the two rectangular bases. SA = p*h + 2+B where B is the area of a rectangular base. 148 = (5 + 5 + 4 + 4)*h + 2*(4*5) @dopegyal_ h = ?

OpenStudy (anonymous):

6?

OpenStudy (anonymous):

148cm^2 surface area. subtract 20cm for top and 20cm for bottom 108cm left for height, 18 liner cm around prism 108/18 for height

OpenStudy (anonymous):

But direct's looks more fancy

OpenStudy (anonymous):

6 it s

Directrix (directrix):

Not all that fancy. Think about a stack of circles, kinda like the toy Slinky. Each circle has circumference 2 pi r and the stack is h tall. So 2 pi r h gives the lateral area. Then, add on the two circular bases. That's the only way I can remember the formula. @ehuman

OpenStudy (anonymous):

I love learning